OpenStudy (anonymous):

What is an equation in point-slope form of the line that passes through the points (7, 5) and (-4, -1) ?

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

pooja195 (pooja195):

Before that we need to find the slope \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (anonymous):

Right....

OpenStudy (anonymous):

@salemlover352

OpenStudy (anonymous):

I am sorry I need all the help I can get.

pooja195 (pooja195):

Im willing to help :) \[\LARGE \frac{-1-5}{-4-7}=slope\]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

-6/-11

OpenStudy (anonymous):

Thats the slope right?

pooja195 (pooja195):

Correct now user this \[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (anonymous):

Oh snap this si where I get stuck

pooja195 (pooja195):

What don't you get?

OpenStudy (anonymous):

All of it, like what to plug in

pooja195 (pooja195):

\[\huge~\rm~\bf y-y_1=m(x-x_1)\]where m=slope=-6/-11 (x1,y1) x1=7 y1=5 (7,5) <---they were the first given pairs of coordinates Now plug everything in and tell me what you get! :) i know you can do it ^-^

OpenStudy (anonymous):

Okay so y - 5 = -6/-11(x - 7)?

pooja195 (pooja195):

perfect! :D

OpenStudy (anonymous):

But thats not even an option though :(

pooja195 (pooja195):

What are your options?

OpenStudy (anonymous):

A. y + 1 = 6/7 (x + 4) B. y + 1 = 6/11 (x + 4) C. y + 4 = 11/6 (x + 1) D. y - 1 = 6/11 (x + 4)

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

2 negatives make a positive slope forgot about that sorry!

pooja195 (pooja195):

And they used the secound pair of coordinates instead of the first

OpenStudy (anonymous):

Oh okay so it would be C?

OpenStudy (anonymous):

I mean A?

pooja195 (pooja195):

\[\huge~\rm~\bf y-y_1=m(x-x_1)\]where m=slope=6/11 (x1,y1) x1=-4 y1=-1 (-4,-1) <---they were the first given pairs of coordinates

OpenStudy (anonymous):

Agh! I meant B and okay

pooja195 (pooja195):

Good :)

OpenStudy (anonymous):

y + 1 = 6/11 (x + 40

OpenStudy (anonymous):

Thanks! SO much, could you help me with 4 more? @pooja195

pooja195 (pooja195):

Ok :)

OpenStudy (anonymous):

Thanks! Here's number 2.

OpenStudy (anonymous):

Which is an equation of the given line in standard form? A. –8x + 9y = 23 B. –8x + 9y = –23 C. –8x + 7y = 25 D. –9x + 8y = –23

OpenStudy (anonymous):

The picture shows the information :). YOu are a lifesavor haha.

pooja195 (pooja195):

Im not good with graph equations @Photon336 can you please help?

OpenStudy (anonymous):

@Photon336 I could really use your help lol.

OpenStudy (photon336):

wait with which one? the one you guys just did?

pooja195 (pooja195):

The last one that was posted

OpenStudy (anonymous):

Yes

OpenStudy (photon336):

@Abbster2015 you can kind of build up the equation by looking at the graph. there are a few hints. \[y = m(x)+b \] the first thing you can do here find the slope of those two points (-4,-1) and (0.5,3) you can do this by using this formula do you know how to do this? \[m = \frac{ y_{2}-y_{1} }{ x_{2}-x_{1} }\]

OpenStudy (anonymous):

Kinda, but I get confused which numbers go in which slots, could you help me with that?

OpenStudy (photon336):

yeah sure, you have to chose which one is y2 and which one is x2 @abbster2015 you can chose any point and designate it as x2 and y2 let's say we chose the point (1/2,3) as x2 and y2  follow?